OpenStudy (anonymous):

2x - y = 10 y = -x - 1 Solve the system of equations (9, 8) correct? (3, -4) no solution

OpenStudy (anonymous):

you're correct

OpenStudy (anonymous):

thanks

OpenStudy (whpalmer4):

No, that is NOT correct. \[2x-y=10\]\[2(9)-8 = 10\]Works \[y=-x-1\]\[8=-9-1\]\[8=-10\]Does not work Must satisfy ALL equations in the system to be a valid solution. rearrange second equation:\[y=-x-1\]\[x+y=-1\] \[2x-y=10\]\[x+y=-1\]Add equations together \[2x+x + y - y = 10 -1\]\[3x + 0y = 9\]\[x = 3\]\[x+y=-1\]\[3+y=-1\]\[y = -4\] Testing: \[2x-y=10\]\[2(3)-(-4) = 10\]\[6+4=10\checkmark\] \[y=-x-1\]\[(-4)=-(3)-1\]\[-4=-4\checkmark\] Therefore \((3,-4\)) is the solution. Alternative solution by substitution: \[2x-y=10\]\[y=-x-1\]Substitute \((-x-1)\) everywhere you see \(y\) in the first equation\[2x-(-x-1) = 10\]\[2x+x + 1 = 10\]\[3x+1=10\]\[3x=9\]\[x=3\]\[y=-x-1 = -(3)-1 = -4\]\((3,-4)\) is solution (not checking it again because I already che